x1 + cost y = 4 + sint (You are welcome to use your calculator for this one.) ОА ов. O C. OD.
Answer #1

Griven that zl= 1 + cost - 4+sinet at - lost 9-4 = simit we know sino & Losht = 1 (al-11²+ (9-4)=1 it is equation of circle c